OpenStudy (jusaquikie):

a chemist runs the balanced chemical reaction shown below by mixing 6.18g o Ca3P2 with 9.49g of H2O, how much of PH3 will be formed asuming 100% yield?

OpenStudy (jusaquikie):

\[Ca _{3}P _{2}(s)+6H _{2}O(l)\rightarrow 3 Ca(OH)_{s}(s)+2 PH _{3}(g)\]

OpenStudy (australopithecus):

http://openstudy.com/study#/updates/50c27afae4b066f22e104a93 I explain limiting reagent

OpenStudy (australopithecus):

If there is a 100% yield you would just be determining the theoretical yield, scroll down to the bottom post of that question

OpenStudy (jusaquikie):

i took the grams of Ca3P2/MM and H20 and got the limiting agent as the CaP2 and that is a 1:1 ratio with the 2PH3

OpenStudy (jusaquikie):

once i figure out the ratio is 1:1 then i take the moles of Ca3P2 then multiply by the MM of PH3 to get the grams right?

OpenStudy (australopithecus):

The ratio between Ca3P2 and PH3 isn't 1:1

OpenStudy (jusaquikie):

i thought i compare the P? P2 in the first to 2P in the second?

OpenStudy (australopithecus):

for every 1 Ca3P2 you produce 2 PH3

OpenStudy (australopithecus):

Show me the moles of Ca3P2 and H2O

OpenStudy (jusaquikie):

\[\frac{ 6.18 }{ 182.18 }=.033922\] \[\frac{ 9.49 }{ 18 }=.52722\] \[\frac{ .52722 }{ .033922 }=15.542\] 15>6 so Ca3P2 is the limiting reactant right?

OpenStudy (australopithecus):

The way I do it is usually convert moles of reagent to product then see which gives me the least product (0.52722/6)2 = (0.033922/1)2 =

OpenStudy (australopithecus):

doing it this way also gives you the moles of product you have so its not like it is a huge hassle

OpenStudy (jusaquikie):

what's the *2 for?

OpenStudy (australopithecus):

you produce 2 PH3

OpenStudy (jusaquikie):

ok so i should have (.03392)*2 moles of PH3

OpenStudy (australopithecus):

yes

OpenStudy (jusaquikie):

then i multiply that by the MM of PH3 and i get the right answer. Thank you very much for your time it's clearer now.

OpenStudy (australopithecus):

yup